The mass of a typical car is

m = 1,500 kg,

so its weight on the Earth's surface is

mg = 14,700 N.

Suppose you have two containers, one with N electrons and another with N protons. These two containers are placed a distance 13 m apart, and the electric force between them is equal to the weight of the car. Find N.
Answer #1

Solution :-

Given,

weight of car , W = 14600N

Mass (m) = 1500kg

let the charge of electrons and protons is q

for force between electrons and protons = weight of car

k * q^2/d^2 = W

9*10^9 * q^2/13^2 = 14700

solving for q

q = 0.0161 C

number of electrons , N = total charge /charge on 1 electron

N = 0.0161/1.602 *10^-19

N = 1.0018 *10^17

the number of electrons and protons is 1.0018 *10^17